OpenStudy (anonymous):

A password question. A password is to consist of 10 alphanumeric characters and must include at least one lower case letter, at least one upper case letter, and at least one digit. (a) How many such passwords are possible? (b) How many are possible if the password must contain at least one lower case letter, at least one upper case letter, and at least one digit which is immediately preceded by a letter?

OpenStudy (anonymous):

I guess I solved the first one. 62^10 all possibility- ( no number+ no U letter+ no L letter)

OpenStudy (anonymous):

(no numbers= L+U, no U= L+N, no L= N+U)

OpenStudy (anonymous):

so answer seems 62^10-(52^10+36^10+30^10)

OpenStudy (anonymous):

@amistre64